1.  A friend promises to pay you RM 700 two years from now if you loan him RM 550 today. What annual interest is your friend offering?   
  2. Calculate the present value of RM 1,000 to be received at the end of 8 years. Assume an interest rate of 7 percent   
  3. Mr A has deposited RM 33,000 today in an account which will earn 10 percent annually. He plans to leave the funds in this account for seven years earning interest. If the goal of this deposit is to cover a future obligation of RM65,000, what recommendation would you make to Mr A? 
  4. Company A is preparing a five-year plan. Today, sales are RM 1,000,000. If the growth rate in sales is projected to be 10 percent over the next five years, what will the RM amount of sales be in year five?
  5. How much would you be willing to pay today for an investment that would return RM 800 each year at the end of each of the next 6 years? Assume a discount rate of 5 percent.
1
Expert's answer
2021-12-07T02:26:03-0500

1.      A = P(1+r/100)n

700 = 550 (1+r/100)2

=16.51%


2.      PV = FV      1

              (1+r)n


PV = 1000 X       1

                                (1+7)

=RM 582


3.      FV = PV (1+r) n

33000 (1+10/100)7

=RM 64,307.66

Mr. a should not deposit the funds in this account as the future value will not fully cover the obligation.


4.      FV = PV (1+r) n

1,000,000 (1+10/100)5

=RM 1,610,510


5.      PV = Payment (1-(1+rn))/r

800(1-(1/(1.056))/0.05

=RM 4060.55